OpenStudy (somegirl1999):

Describe how you would estimate the square root of a number that is not a perfect square without using a calculator.

OpenStudy (somegirl1999):

i need the answer lol

OpenStudy (anonymous):

well what if you found perfect squares that this non perfect number is inbetween

OpenStudy (somegirl1999):

I would .... I'm not sure I'm not good at math.

OpenStudy (anonymous):

alright so say you have a number 7...

OpenStudy (anonymous):

so \[\sqrt{7}\]it is not the perfect square however 4 and 9 are

OpenStudy (anonymous):

and we know that it's square root is greater than square root for 4 and less than 9

OpenStudy (anonymous):

\[\sqrt{4}<\sqrt{7}<\sqrt{9}\] \[2<\sqrt{7}<3\]

OpenStudy (somegirl1999):

thank you

OpenStudy (anonymous):

so now we know that it's between 2 and 3, take the number that it between the two outer numbers 9-4=5 ... use this as a partitions in that difference between 2 and 3 (1). This means that each partition is .2 (1/5). since there is a difference of 3 additions of 1, you'd add 3 .2's to the number

OpenStudy (anonymous):

so my estimate would be \[\sqrt{7}=2.6\]

OpenStudy (anonymous):

the actual is \[sqrt{7}=2.64\]

OpenStudy (somegirl1999):

Thank you so much